# Comparing Cluster-Based Cross-Validation Strategies for Machine Learning Model Evaluation

### v1

@@ II-D K-Fold Cross-Validation
Among cross-validation techniques, the most commonly used is K-Fold Cross-Validation. In this method, the dataset is divided into $K$ parts, called folds. Then, the model is trained on $K-1$ parts and tested on the remaining part. This process is repeated $K$ times, with a different fold being used for validation in each iteration. Finally, the model’s performance is computed as the arithmetic mean of the performance across all iterations. Figure 2 illustrates the training and testing process using K-Fold Cross-Validation with five folds.
@@ III-B Distribution Optimally Stratified Cross-Validation (DOBSCV)
DOBSCV works similarly to DBSCV, starting with a random instance from the dataset. However, instead of finding the nearest instance of the same class, it identifies the $k-1$ nearest neighbors (where $k$ is the number of folds) of the current instance from the same class and assigns each one to a different fold. This process is repeated independently for each class, similarly to DBSCV, until all instances have been assigned.
@@ IV-A Datasets
Table I presents the list of datasets used in the experiments, divided into balanced and imbalanced categories. This division was made to analyze the effectiveness of cross-validation techniques in distinct class distribution scenarios, using appropriate metrics for each type. The imbalance of the PMLB datasets is measured by computing the sum of the squared distances between the proportion of instances in each class and the perfectly balanced distribution in the dataset, as shown in Equation 1, where $n_{i}$ is the number of instances of class $i$, $K$ is the total number of classes, and $N$ is the dataset size.
@@
[…]
@@
iris 150 4 3 0,00 4
analcatdata_germangss 400 5 4 0,00 4
movement_libras 360 90 15 0,00 5
sonar 208 60 2 0,00 4
vowel 990 13 11 0,00 4
@@ IV-B Classification Algorithms and Hyperparameter Optimization
The optimized hyperparameters are shown in Table II. For LR and SVM, the hyperparameter $C$, which controls model regularization, was tuned, with lower values imposing stronger regularization. For SVM, the $gamma$ parameter was also adjusted, which defines the influence of each data point in the decision function. For DT and RF, the $max_{d}epth$ parameter was tuned to limit the depth of the trees, controlling model complexity and preventing overfitting. The tested values were selected to cover a wide range of scenarios, ensuring that both simple and complex models were evaluated, and enabling the selection of the most suitable hyperparameters according to the specific characteristics of each dataset.
@@ IV-C Clustering Algorithms and Hyperparameter Definition
The clustering algorithms used in the cross-validation procedures during the experiments require hyperparameters to be provided as input. While K-Means and Agglomerative Clustering require the number of clusters, DBSCAN depends on the parameters $\varepsilon$ and $min\_samples$.
@@
[…]
In the SCBCV Mini and KCBCV Mini algorithms, which employ Mini-Batch K-Means, the hyperparameter $batch_{s}ize$ was set to the default value of 102410241024. The batch size determines the number of samples processed in each iteration of the algorithm, directly affecting both computational efficiency and clustering accuracy. A larger batch size allows more data to be processed at once, which can accelerate the algorithm’s convergence, but also requires more memory and may result in reduced accuracy for smaller datasets.
@@ IV-D Class-Stratified K-Means Cross-Validation
In its input parameters, SCBCV uses the dataset $X$, containing the features to be analyzed, and $y$, which represents the corresponding class labels. The parameter $k_{s}plits$ defines the number of folds for cross-validation, while $k_{c}lusters$ specifies the number of clusters to be formed by the K-Means algorithm. The $rng$ parameter ensures experiment reproducibility through a random number generator, and $minibatch_{k}means$ indicates whether the optimized Mini-Batch K-Means version will be used to efficiently handle large volumes of data.
@@
[…]
@@
3:  if $minibatch\_kmeans$ then
[…]
@@
[…]
@@
10:   $class\_dist\leftarrow$ executa $kmeans$ e retorna matriz de distâncias de cada ponto a cada centróide
11:   $clusters\_index\leftarrow$ obtém índices dos clusters de cada elemento em $class\_dist$
12:   $clusters\leftarrow$ listas vazias para cada um dos $k\_clusters$
13:   for index em $class\_dist$ do
@@
22:  $folds\leftarrow$ distribui $index\_list$ entre $k\_splits$ $folds$ de forma circular
23:  return $folds$
@@ IV-E Evaluation Metrics
[…]

### v2

@@ II-D K-Fold Cross-Validation
Among cross-validation techniques, the most commonly used is K-Fold Cross-Validation. In this method, the dataset is divided into $K$ parts, called folds. Then, the model is trained on K1K-1 parts and tested on the remaining part. This process is repeated $K$ times, with a different fold being used for validation in each iteration. Finally, the model’s performance is computed as the arithmetic mean of the performance across all iterations. Figure 2 illustrates the training and testing process using K-Fold Cross-Validation with five folds.
@@ III-B Distribution Optimally Stratified Cross-Validation (DOBSCV)
DOBSCV works similarly to DBSCV, starting with a random instance from the dataset. However, instead of finding the nearest instance of the same class, it identifies the k1k-1 nearest neighbors (where $k$ is the number of folds) of the current instance from the same class and assigns each one to a different fold. This process is repeated independently for each class, similarly to DBSCV, until all instances have been assigned.
@@ IV-A Datasets
Table I presents the list of datasets used in the experiments, divided into balanced and imbalanced categories. This division was made to analyze the effectiveness of cross-validation techniques in distinct class distribution scenarios, using appropriate metrics for each type. The imbalance of the PMLB datasets is measured by computing the sum of the squared distances between the proportion of instances in each class and the perfectly balanced distribution in the dataset, as shown in Equation 1, where $n_{i}$ is the number of instances of class $i$, $K$ is the total number of classes, and NN is the dataset size.
@@
The higher the value of index II, the greater the class imbalance, with II approaching 1 when almost all instances belong to a single class. In this study, datasets with an imbalance index greater than $0,20$ were classified as imbalanced [3]. As an example, the haberman dataset used in the experiment has an imbalance index of $0.22$ with two classes, where 26.5% of the data belongs to one class and $73.5\%$ to the other, characterizing a considerable imbalance. Another dataset used, dis, presents an imbalance index of $0.94$, also with two classes, where $98.5\%$ of the instances belong to one class and only $1.5\%$ to the other, illustrating an extreme and significant imbalance.
@@
iris 150 4 3 0,$0$ 4
analcatdata_germangss 400 5 4 0,$0$ 4
movement_libras 360 90 15 0,$0$ 5
sonar 208 60 2 0,$0$ 4
vowel 990 13 11 0,$0$ 4
@@ IV-B Classification Algorithms and Hyperparameter Optimization
The optimized hyperparameters are shown in Table II. For LR and SVM, the hyperparameter $C$, which controls model regularization, was tuned, with lower values imposing stronger regularization. For SVM, the gammagamma parameter was also adjusted, which defines the influence of each data point in the decision function. For DT and RF, the $max_{d}epth$ parameter was tuned to limit the depth of the trees, controlling model complexity and preventing overfitting. The tested values were selected to cover a wide range of scenarios, ensuring that both simple and complex models were evaluated, and enabling the selection of the most suitable hyperparameters according to the specific characteristics of each dataset.
@@ IV-C Clustering Algorithms and Hyperparameter Definition
The clustering algorithms used in the cross-validation procedures during the experiments require hyperparameters to be provided as input. While K-Means and Agglomerative Clustering require the number of clusters, DBSCAN depends on the parameters $\varepsilon$ and min_samplesmin\_samples.
@@
For the DBSCAN algorithm, the procedure suggested by Sander et al. [4] was followed. The $min_{s}amples$ parameter, which defines the minimum number of neighboring points required to form a cluster, was set as $2*\text{number of columns}$ in the dataset. The $\varepsilon$ parameter, which defines the radius within which data points are considered neighbors, was calculated using a $k$-distance graph, where the distance to the $k$-th nearest neighbor, with k = min_samples - 1, is plotted in descending order. The appropriate value for $\varepsilon$ corresponds to the point at which a significant change in the slope of the graph occurs.
In the SCBCV Mini and KCBCV Mini algorithms, which employ Mini-Batch K-Means, the hyperparameter $batch_{s}ize$ was set to the default value of $1024$. The batch size determines the number of samples processed in each iteration of the algorithm, directly affecting both computational efficiency and clustering accuracy. A larger batch size allows more data to be processed at once, which can accelerate the algorithm’s convergence, but also requires more memory and may result in reduced accuracy for smaller datasets.
@@ IV-D Class-Stratified K-Means Cross-Validation
In its input parameters, SCBCV uses the dataset $X$, containing the features to be analyzed, and $y$, which represents the corresponding class labels. The parameter $k_{s}plits$ defines the number of folds for cross-validation, while $k_{c}lusters$ specifies the number of clusters to be formed by the K-Means algorithm. The rngrng parameter ensures experiment reproducibility through a random number generator, and $minibatch_{k}means$ indicates whether the optimized Mini-Batch K-Means version will be used to efficiently handle large volumes of data.
@@
1:procedure SCBCV($X$, $y$, k_splitsk\_splits, k_clustersk\_clusters, rngrng, minibatch_kmeansminibatch\_kmeans)
@@
3:  if minibatch_kmeansminibatch\_kmeans then
4:   $kmeans\leftarrow$ MiniBatchKMeans(k_clustersk\_clusters, rngrng)
@@
6:   $kmeans\leftarrow$ KMeans(k_clustersk\_clusters, rngrng)
@@
10:   $class\_dist\leftarrow$ executa kmeanskmeans e retorna matriz de distâncias de cada ponto a cada centróide
11:   $clusters\_index\leftarrow$ obtém índices dos clusters de cada elemento em class_distclass\_dist
12:   $clusters\leftarrow$ listas vazias para cada um dos k_clustersk\_clusters
13:   for index em class_distclass\_dist do
@@
22:  $folds\leftarrow$ distribui index_listindex\_list entre k_splitsk\_splits foldsfolds de forma circular
23:  return foldsfolds
@@ IV-E Evaluation Metrics
$\text{F1-score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)